OpenStudy (nudeoua):

Evaluate [(40 + 5) − 3^2] ÷ 9*2 8.6 43 8 2 I got two, but the right answer was eight.. This is what I did : {(40+5)-3^2}/9*2 {45-9}/9*2 36/18 2 Where did I go wrong?

OpenStudy (aaronandyson):

{45-9}/9*2<-- till this step everything is correct.

OpenStudy (aaronandyson):

45 - 9 = ?

OpenStudy (aaronandyson):

36 right?

OpenStudy (aaronandyson):

Now, \[\frac{ 36 }{ 9 } \times 2\]

OpenStudy (aaronandyson):

36 divide by 9 gives 4

OpenStudy (aaronandyson):

and 4*2 = 8

OpenStudy (aaronandyson):

Getting it @nudeoua ??

OpenStudy (nudeoua):

I got it, thanks!